OpenStudy (konradzuse):

Classify the graph of the eq 9x^2+25y^2-54x+250y+481=0

OpenStudy (konradzuse):

I know it's an ellipse, but I have to turn it into the proper form. What I want to know is how do I find h and k, or what shall I do?

OpenStudy (konradzuse):

@asnaseer @TuringTest @lgbasallote :)

OpenStudy (lgbasallote):

have you tried completing the square?

OpenStudy (konradzuse):

yup

OpenStudy (konradzuse):

I tried to split it up into (9x^2 - 54x) and (25y^2+250y+500)-19 and a few others, but couldn't complete the square with y...

OpenStudy (konradzuse):

I know viper, I need to prove that it's an ellipse, so I need to get it into proper form.

OpenStudy (lgbasallote):

9x^2+25y^2-54x+250y+481=0 transpose 481 9x^2 + 25y^2 -54x + 250y = -481 now do completing the square 9x^2 - 54x + 25y^2 + 250y = -481 9(x^2 - 6x) + 25(y^2 + 10y) = -481 9(x^2 - 6x + 9) + 25(y^2 + 10y + 25) = -481+ 81 + 625 9(x-3)^2 + 25(y + 5)^2 = 225 divide everything by 225... \[\frac{(x-3)^2}{25} + \frac{(y+5)^2}{9} = 1\] does that help?

OpenStudy (lgbasallote):

@KonradZuse walked out on me again :'C

OpenStudy (konradzuse):

Dude this site keeps crashing on me haha...

OpenStudy (konradzuse):

It just logged me out :(.

OpenStudy (lgbasallote):

haha lol

OpenStudy (konradzuse):

interesting how you solved this, thanks iggy :).

OpenStudy (konradzuse):

ok I got another one that's a hyperbola... x^2-4y^2-2x-16y-31=0

OpenStudy (konradzuse):

(X^2-2x) - (4y^2-16y) = 31

OpenStudy (lgbasallote):

+16y

OpenStudy (lgbasallote):

and honestly..i had no idea what i was doing...im weak in graphs....it was just a fluke..i was hoping it would come out right in the end =)))

OpenStudy (konradzuse):

(x^2-2x + 1) - 4(y^2+4y +4) = 31 -1 -16

OpenStudy (konradzuse):

LOL dude don't tell me that :(

OpenStudy (lgbasallote):

should be 31+1

OpenStudy (konradzuse):

@TuringTest what is he doijng to me?

OpenStudy (konradzuse):

+1? Isn't it +1 on the other side?

OpenStudy (lgbasallote):

lol it was right

OpenStudy (konradzuse):

oh wait derp.

OpenStudy (lgbasallote):

lol what are you saying

OpenStudy (konradzuse):

+16 also then....

OpenStudy (lgbasallote):

nope

OpenStudy (lgbasallote):

it's -16

OpenStudy (lgbasallote):

-4 x 4 = -16

OpenStudy (konradzuse):

cuz of the neg in front?

OpenStudy (konradzuse):

I gues ssooo..

OpenStudy (lgbasallote):

wiat....

OpenStudy (konradzuse):

so it's now (x-1)^2 -4(y+2) = 16

OpenStudy (lgbasallote):

something smells wrong here

OpenStudy (konradzuse):

I blamr you dude...

OpenStudy (lgbasallote):

lol

OpenStudy (lgbasallote):

i did the ellipse right :p

OpenStudy (konradzuse):

:)

OpenStudy (lgbasallote):

anyway i'll be right back..i'll help if it wasnt solved by the time i get back

OpenStudy (konradzuse):

isn't it just going to be (x-1)^2/16 - (y+2)^2/4 = 1?

OpenStudy (konradzuse):

look the same to me, thanks :)

OpenStudy (lgbasallote):

lol yeah...guess i didnt see that one...good one @KonradZuse

OpenStudy (konradzuse):

yeah iggy!!!! :)